The volume of hydrogen sulphide it stp when 1.5g of sodium hydroxide react with hydrogen sulphide.


(NA=23,H=1,S=32,O=16)



1
Expert's answer
2022-06-03T04:17:03-0400

Solution:

The molar mass of sodium hydroxide (NaOH) is (23 + 16 + 1) = 40 g/mol

Therefore,

Moles of NaOH = (1.5 g NaOH) × (1 mol NaOH / 40 g NaOH) = 0.0375 mol NaOH


Balanced chemical equation:

2NaOH + H2S → Na2S + H2O

According to stoichiometry:

2 mol of NaOH react with 1 mol of H2S

Thus 0.0375 mol of NaOH react with:

(0.0375 mol NaOH) × (1 mol H2S / 2 mol NaOH) = 0.001875 mol H2S


At STP, one mole of any gas occupies a volume of 22.4 L

Thus 0.001875 mol of H2S occupy:

(0.001875 mol H2S) × (22.4 L H2S / 1 mol H2S) = 0.42 L H2S


Answer: The volume of hydrogen sulphide (H2S) is 0.42 L